OpenStudy (anonymous):

Limit Evaluation:

OpenStudy (anonymous):

\[\Huge \lim_{x \rightarrow 0} \frac{(2+h)^3 -8}{h} \]

OpenStudy (experimentx):

it's a function of h and x->0

OpenStudy (experimentx):

if it's h->0, expand that cube ... cancel out 8, then cancel out h on top and bottom ad get your answer

OpenStudy (anonymous):

How do I do that?

OpenStudy (experimentx):

(2+h)^3 = 8 + h^3+3*2*h(h+2)

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=%282%2Bh%29^3-8

OpenStudy (anonymous):

or use L' hospital's rule..ans=4

OpenStudy (experimentx):

my bad (2+h)^3 = 8 + h^3+3*h(h+2)

OpenStudy (experimentx):

must be 6

OpenStudy (anonymous):

but by differenting numerator becomes 3*(h+2)^2 = 12

OpenStudy (experimentx):

Oh ... so .. (2+h)^3 = 8 + h^3+3*2*h(h+2) was right ... lol